The footballer kicked the ball and it was caught by the opposing goalkeeper, 4 metres above the ground. The ball was traveling at 10 m/s^2.Calculate the kinetic energy of the ball just before it was caught.


1
Expert's answer
2022-03-15T10:36:23-0400

Solution.

"v_0=10m\/s;"

"h=4m;"

"W_k-?;"

"\dfrac{mv_0^2}{2}=mgh_{max}\implies h_{max}=\dfrac{v_0^2}{2g};"

"h_{max}=\dfrac{100}{2\sdot9.81}=5.1m;"

"mgh_{max}=mgh+\dfrac{mv^2}{2}\implies W_k= \dfrac{mv^2}{2}=mgh_{max}-mgh;"

"W_k=m(gh_{max}-gh);"

"W_k=m(9.81\sdot5.1-9.81\sdot4)=10.79mJ;"

Answer: "W_k=10.79m J."